Grade 12Physical Chemistry

a gas is found to have a formula [CO]x. its vapour density is 70, the x is
a)3 b)3.5 c)5 d) 6.5

Dhwani shah
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
"Mol. Wt. =2*vapour density
hence, Mol Wt. Of carbon =12u
Mol Wt. Of O=16u
Mol Wt. Of (CO) x =(12+16)x =28x
Now, According to Formula
28x =2*70
28x =140
x=5 is possible value
Molecular mass is equal to 2 X vapour density 2 X 70 is equal to x times atomic mass of carbon + atomic mass of oxygen
(12 + 16 )x = 28 x show 28 x = 140 and x = 5